OpenStudy (anonymous):

Use the triangle at the right. Find the length of the missing side. Show your work Did i do the first one right? I don't want to do the second one wrong too. 1.a = 16, b = 63 2. b = 2.1, c = 2.9 16^2 =256, 63^2=3,969 256+3969 = 4225 now we need to find the square root by separating the 4225 into two separate numbers square root of 42 is 6 now find the largest root in 42. Which is 36. 6*6 = 36 and the square of 25 is 5 add the roots This give me the answer for c length. 6 and 5 is 65. a=256 b=3969 and c = 65

OpenStudy (solomonzelman):

Do you have a picture or other description of the triangle you are dealing with in problem 1 and/or problem 2?

OpenStudy (anonymous):

yup you need a picture

OpenStudy (anonymous):

The picture probably wouldn't help it's just a right triangle but okay

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

this picture will definitely help, because I would have thought that b is the hypotenuse

OpenStudy (solomonzelman):

i mean in question 1 i would have thought so, but this picture gives me that c=hypotenuse in both cases, and I know it is a right triangle

OpenStudy (anonymous):

Alright.

OpenStudy (solomonzelman):

I want to correct you though. The set up is a little different, that is 1. Here is the rule. A right triangle with sides a, b, c (where the hypotenuse is side c), must satisfy the following statement: a²+b²=c²

OpenStudy (solomonzelman):

this statement is known as the pythagorean theorem.

OpenStudy (solomonzelman):

Now, you are given your two smaller legs a and b are 16 and 63 (respectively). And you are missing the hypotenuse, so this is what you would do. (16)² + (63)² = c²

OpenStudy (solomonzelman):

then you simplify the left hand side, and solve for c (just by talking the square root of both sides)

OpenStudy (anonymous):

So i had to square c?

OpenStudy (anonymous):

Take the square root of a and b?

OpenStudy (solomonzelman):

(16)² + (63)² = c² 4225 = c² then do this: \(\sqrt{4225}\) = \(\sqrt{{\rm c}^2}\)

OpenStudy (solomonzelman):

this way you are able to solve for c. (normally the square root will give you the ±, as you know, but in this case since distance or sidelength can not be negative, you diregard any negative solutions)

OpenStudy (anonymous):

Alright. is c 65? @SolomonZelman

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

Did i do the problem right?

OpenStudy (solomonzelman):

i wrote it all out and still doesn't take that much space and time..... ok, now question 2...

OpenStudy (anonymous):

Kay.

OpenStudy (solomonzelman):

oh, I deleted that.... ------------------ a²+b²=c² 16²+63²=c² 4225=c² √4225 = √c² c=65 ------------ reposted it.

OpenStudy (solomonzelman):

ok, in question 2, you are given c=2.9 b=2.1

OpenStudy (solomonzelman):

can you plug in this information into the a²+b²=c² ?

OpenStudy (anonymous):

Sure a^2+ b^2=c^2

OpenStudy (solomonzelman):

but, you are given the c and b, so you can g ahead and plug in 2.9 for c, and 2.1 for b.

OpenStudy (anonymous):

2.9^2 + 2.1^2

OpenStudy (solomonzelman):

no,

OpenStudy (solomonzelman):

The pythagorean theorem is: a²+b²=c² where c is the hypotenuse and a & b are two legs. you are given that your c (which is hypotenuse as well) is 2.9 and you are given that your b is 2.1 (your missing side is a)

OpenStudy (anonymous):

Alright

OpenStudy (solomonzelman):

please take a shot to plug in your values (into a²+b²=c²)

OpenStudy (anonymous):

2.9^+2.1=c^2

OpenStudy (solomonzelman):

your c is given, but a is not

OpenStudy (solomonzelman):

it is like this: a² + (2.1)² = (2.9)²

OpenStudy (anonymous):

Oh sorry.

OpenStudy (solomonzelman):

it's ok...

OpenStudy (anonymous):

2.9^+2.1=c^2

OpenStudy (solomonzelman):

a² + (2.1)² = (2.9)²

OpenStudy (anonymous):

Wifi is bad okay a^2 +(2.1)^2 =( 2.9)^2

OpenStudy (solomonzelman):

because you are given: c=2.9 b=2.1 so, the missing side is a. Our theorem is: a² + b² = c² so lets plug in everything plugging 2.9 for c plugging 2.1 for b you get: a² + (2.1)² = (2.9)²

OpenStudy (solomonzelman):

ok now solve for a

OpenStudy (anonymous):

Alright one moment

OpenStudy (anonymous):

Do i add the c value or divide or do it the same thing i did in my last problem.

OpenStudy (anonymous):

@SolomonZelman you there?

OpenStudy (solomonzelman):

you first calculate the values of (2.1)² and (2.9)²

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

4.41= 8.41. Do i subtract next?

OpenStudy (solomonzelman):

a² + (2.1)² = (2.9)² without a calculator: 21•20=420 21•21=420+21=441 so 2.1² = 4.41 30•30=900 30•29=900-30=870 29•29=870-29=841 so 2.9²=8.41

OpenStudy (solomonzelman):

just demonstrating another technique.

OpenStudy (solomonzelman):

anyway a² + (2.1)² = (2.9)² a² + 4.41 = 8.41 yes you subtract 4.41 from both sides

OpenStudy (anonymous):

Okay.

OpenStudy (solomonzelman):

a² + 4.41 \(\small \color{red}{-4.41}\)= 8.41\(\small \color{red}{-4.41}\)

OpenStudy (solomonzelman):

and from this u get?

OpenStudy (anonymous):

4.00

OpenStudy (solomonzelman):

yes, or just 4 :)

OpenStudy (solomonzelman):

so, a²=4 correct?

OpenStudy (anonymous):

Yeah it looks correct i don't think it can be factored anymore

OpenStudy (solomonzelman):

no, there is no factoring here:) so we got a²=4 what do you think your next (and final) step is?

OpenStudy (anonymous):

putting them together like 2.1^2+2.9^2= 4^2

OpenStudy (solomonzelman):

u just take the square root of both sides \(a^2=4\) \(\color{red}{\sqrt{\color{black}{a^2}}}=\color{red}{\sqrt{\color{black}{4} }}\)

OpenStudy (solomonzelman):

a = ?

OpenStudy (anonymous):

4^2?

OpenStudy (solomonzelman):

a square root of a 4 is?

OpenStudy (anonymous):

2?

OpenStudy (solomonzelman):

yes so a=2

OpenStudy (anonymous):

Alright is there anything else?

OpenStudy (solomonzelman):

no, you found the missing side in both of the problems.

OpenStudy (anonymous):

Alright thanks for your help :)

OpenStudy (solomonzelman):

yw